OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through the points (-2, 7) and (3, 12) Almost done with my practice problems?

OpenStudy (anonymous):

slope intercept form of this line is: y = x+9 http://www.hackmath.net/en/calculator/line-slope?x0=-2&y0=7&x1=3&y1=12&submit=Solve y = kx + q Manually calculate slope k = dy / dx = (12-7)/(3+2) = 1 and for q = y - x = 7-(-2) = 9 Thats easy

OpenStudy (anonymous):

AWESOME C: